A 4.60-mu F capacitor that is initially uncharged
Answer #1

Capacity = C = 4.60 uF

Resistance = R = 7.50 kΩ = 7500 Ω   

emf of source = E=245 V negligible internal resistance.

A)Just after the circuit is completed, the voltage drop across the capacitor is zero

Capacitor acts like a straight wire, thus V = 0 V

B) Just after the circuit is completed, the voltage drop across the resistor is 245 V

Since no voltage drop occurs across the capacitor, all the voltage drops across the resistor: Vr = 245 V

C) just after the circuit is completed, the charge on the capacitor is Q = 0

D)Just after the circuit is completed, the current through the resistor is

I = (Vapplied - Vcap)/R = 245 / 7500 = 0.0327 A


E) A long time after the circuit is completed the valuesof the quantities in parts (a)-(d) are

A) After long time,the voltage drop across the capacitor is 245 V

After a long time, capacitor acts like a very large resistor (or break). So voltage across the capacitor becomes Vc = 245 V

B) After long time,the voltage drop across the resistor is zero

Since all the voltage drops across the capacitor after a very long time, Vr = 0 V

C) the charge on the capacitor is Q = C*V = 4.60*10-6*245 = 1.127*10-3 coulombs

D) After long time,the current through the resistor is zero

Again, capacitor acts line a break or a giant resistor, using ohm’s law with R = infinity, I = 0 A